[Jenne, Koji]
I was trying to lock the Yarm, and saw that I was not getting signals to go between the LSC and SCY models. I had digital zeros for TRY, and when I overrode the trigger and tried to force signal to ETMY, I had digital zeros at the SUS-ETMY_LSC input. The corresponding filter bank in the rfm model was receiving signals, so the Dolphin connection between LSC and SUS was okay, it was just the RFM connection going to the end station that wasn't succeeding.
Koji restarted the c1scy model, and then went inside the IFO room, and found that the SUS IO chassis power was off. We must have accidentally turned it off while we were in there earlier. Koji turned on the power, and also restarted the rfm model, and we now have real signals going back and forth.
Yarm is locked, ASS worked nicely, etc, etc, so things seem normal again (with the Yarm....ETMX stuff is still out of order). |